OpenStudy (anonymous):

Thirty-two is what percent of 80? A. 25% B. 40% C. 52% D. 250%

OpenStudy (alexandervonhumboldt2):

80*x/100=32

OpenStudy (anonymous):

whoa

OpenStudy (alexandervonhumboldt2):

solve for x

OpenStudy (anonymous):

i just posted this question the 4 people joined

OpenStudy (agl202):

B.

OpenStudy (anonymous):

now im confussed

OpenStudy (agl202):

you turn it into fraction of 32 over 80 then divide 32 by 80 you can turn it into a decimal and that's 0.4. move the decimal and you get 40%

OpenStudy (anonymous):

hey i was going to say that

OpenStudy (anonymous):

b

OpenStudy (michele_laino):

please what is \[80 \cdot \frac{{40}}{{100}} = ...?\]

OpenStudy (anonymous):

40

OpenStudy (michele_laino):

no, I think it is 32, since we have: \[80 \cdot \frac{{40}}{{100}} = \frac{{3200}}{{100}} = 32\]

OpenStudy (abcdefghijklmq):

B. Let x% of 80 = 32 or, x/100*80 = 32 or, x = 32*100/80 So x= 40 Thus 40% of 80 is 32